OpenStudy (anonymous):

What is 6!/2! ? A. 360 B. 3 C. 6 D. 720

OpenStudy (butterfield1215):

B.

OpenStudy (butterflydreamer):

hint: 6! = 6 * 5 * 4 * 3 * 2 * 1= ? 2! = 2*1 = ?

jagr2713 (jagr2713):

@Butterfield1215 It is better to guide the user to an answer rather than doing all the work yourself. We discourage giving out answers without letting the asker at least take some part in the analysis - the aim of this site is to teach. Please refer to our code of conduct at http://openstudy.com/code-of-conduct for more information. Thank you.

OpenStudy (anonymous):

I got it

OpenStudy (butterflydreamer):

what'd you get? :)

OpenStudy (anonymous):

it A

OpenStudy (butterflydreamer):

good job! That's correct

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

your the best

OpenStudy (butterflydreamer):

haha thank you :)

OpenStudy (anonymous):

well now I don't need to use the caucator